How many minimum no of sound sources of 80db loudness need to constructively interfere to produce a loudness of at least 120db?

Grade:11

1 Answers

Riddhish Bhalodia
8 years ago
During constructive interference of two equal amplitude waves(A each) the total amplitude doubles. Hence asI \propto A^2implies intensity increases by 4 times.
Say n waves are required, following the above logic theI_{new} = n^2I.
Thus
80db = klog(I)
120db = klog(n2I)

we know k and hence we can solve for n